Circular No. 9027

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 E1 (ITAGAKI) M. Kusiak, Astronomical Observatory, Jagiellonian University, reports

that this comet is apparently visible in SOHO SWAN ultraviolet images (additional details are given on CBET 1721). Visual total-magnitude and coma-diameter estimates: Mar. 14.85 UT, 9.6, 4′.5 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector; zodiacal light); 15.103, 10:, 3′ (A. Hale, Cloudcroft, NM, U.S.A., 41-cm reflector; clouds; small condensed area within a diffuse coma); 15.95, 9.6, 3′.5 (M. Goiato, Araçatuba, Brazil, 22-cm reflector).

Further to IAUC 9026, additional astrometry has been published by B. G. Marsden on MPEC 2009-E69, together with the following preliminary parabolic orbital elements and an ephemeris.

T = 2009 Apr. 7.385 TT  ω = 47.^{o}130
Ω = 105.780 }2000.0
q = 0.60809 AU  i = 128.398

COMETS C/2008 W2, C/2008 W3, C/2008 W4, C/2008 W5 (SOHO)

Further to IAUC 9024, additional Kreutz sungrazers have been found on SOHO website images, with their “discovery” observations tabulated below. C/2008 W2 was tiny and condensed (mag ∼ 7.5). C/2008 W3 was stellar in appearance (mag ∼ 7–7.5) in C3 images, and quite condensed in C2 images. C/2008 W4 was tiny and stellar in appearance (mag ∼ 8). C/2008 W5 was stellar in appearance (mag ∼ 6.5) in C3 images, and fairly condensed with a very short, diffuse tail in C2 images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 W2 Nov. 16.317 15^{h}25.^{m}5 −20^{o}29′ C2  JR 2009-C47
C/2008 W3  23.388 15 49.0 −23 40 C3/2 AK 2009-C47
C/2008 W4 23.521 15 56.0 −21 11 C2 HS 2009-C56
C/2008 W5  23.888 15 51.2 −23 46 C3/2 JR 2009-C56

DONATIONS TO THE CENTRAL BUREAU The Central Bureau gratefully acknowledges recent receipt of monetary

donations from the following individuals: William Dillon (Missouri City, TX, U.S.A.), William Liller (Viña del Mar, Chile), and Alan Gilmore (Lake Tekapo, New Zealand). Additional donations will greatly help the work of the Central Bureau and are encouraged.

2009 March 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green